(2xz-yz)dx+(2yz-zx)dy-(x2-zx+y2)dz=0


1
Expert's answer
2021-04-13T23:31:14-0400

We have given the differential equation,

"(2xz-yz)dx+(2yz-zx)dy-(x^2-zx+y^2)dz=0"

This equation can be solved using the Pfaffian method,


Firstly we have to check whether the given differential equation is integrable or not.


Compare it with "Pdx+Qdy+Rdz = 0"

We get,

"P = (2xz-yz)"

"Q = (2yz-xz)"

"R = -(x^2-zx+y^2)"


Now, "P(\dfrac{\delta Q}{\delta z}-\dfrac{\delta R}{\delta y})+Q(\dfrac{\delta R}{\delta x}-\dfrac{\delta P}{\delta z})+R(\dfrac{\delta P}{\delta y}-\dfrac{\delta Q}{\delta x}) =0"


After putting values we get "P(\dfrac{\delta Q}{\delta z}-\dfrac{\delta R}{\delta y})+Q(\dfrac{\delta R}{\delta x}-\dfrac{\delta P}{\delta z})+R(\dfrac{\delta P}{\delta y}-\dfrac{\delta Q}{\delta x}) =0"


Hence the given differential equation is integrable.

Now to find the solution of the differential equation consider 'z' as constant then "dz =0"


Hence, "Pdx+Qdy =0"

then,

"(2xz-yz)dx+(2yz-xz)dy =0"

which is a exact differential equation whose solution can be given as


"x^2z-xyz + y^2z = C"

which is of the form "u(x,y,z) = C"


Hence, "u(x,y,z) = x^2z-xyz + y^2z"  

Now, we are introducing two variables "\mu" and "k"

whose values can be calculates as


"\mu = \dfrac{1}{P}\dfrac{\delta u}{\delta x}"


"= \dfrac{1}{(2xz-yz)}{2xz-yz}{}"


"=1"

and, "k = \mu R - \dfrac{\delta u}{\delta z}"


"k = x^2-zx+y^2-x^2-xy+y^2"

"k = -xy-zx"

"k = -x(y+z)"

Solution of differential equation can be find out by solving the differential equation,

"\dfrac{\delta u}{\delta z}+k = C"


"x^2-xy + y^2-x(y+z) = C"


This is the solution of the given differential equation.